# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ Abstract
Motivated by recent work connecting Higgs phases to symmetry protected topological (SPT) phases, we investigate the interplay of gauge redundancy and global symmetry in lattice gauge theories with Higgs fields in the presence of a boundary. The core conceptual point is that a global symmetry associated to a Higgs field, which is pure-gauge in a closed system, acts physically at the boundary under boundary conditions which allow electric flux to escape the system. We demonstrate in both Abelian and non-Abelian models that this symmetry is spontaneously broken in the Higgs regime, implying the presence of gapless edge modes. Starting with the U(1) Abelian Higgs model in 4D, we demonstrate a boundary phase transition in the 3D XY universality class separating the bulk Higgs and confining regimes. Varying the boundary coupling while preserving the symmetries shifts the location of the boundary phase transition. We then consider non-Abelian gauge theories with fundamental and group-valued Higgs matter, and identify the analogous non-Abelian global symmetry acting on the boundary generated by the total color charge. For SU($N$) gauge theory with fundamental Higgs matter we argue for a boundary phase transition in the O($2N$) universality class, verified numerically for $N=2,3$. For group-valued Higgs matter, the boundary theory is a principal chiral model exhibiting chiral symmetry breaking. We further demonstrate this mechanism in theories with higher-form Higgs fields. We show how the higher-form matter symmetry acts at the boundary and can spontaneously break, exhibiting a boundary confinement-deconfinement transition. We also study the electric-magnetic dual theory, demonstrating a dual magnetic defect condensation transition at the boundary. We discuss some implications and extensions of these findings and what they may imply for the relation between Higgs and SPT phases.
I Introduction
@@
The redundancy inherent to gauge theories leads to important subtleties. In particular, while the theories are local, the physical gauge-invariant objects are non-local Wegner-Wilson string loops [1, 2]. This implies a tension in describing the spontaneous breaking of symmetries in the presence of dynamical gauge fields. Whereas many condensed matter systems demonstrate symmetry lowering phase transitions governed by spontaneous breaking of a global symmetry, gauge redundancy, being unphysical, cannot be broken. This fact, enshrined in Elitzur’s theorem, belies a rich landscape of different phases separated by phase transitions whose study began systematically in the 1970s. While Landau theory successfully accounts for a broad range of phase transitions in correlated many-body systems, the order parameters for theories with dynamical gauge fields are generically non-local, raising the question of how to understand the nature of the phase transitions in such theories. Recent advances generalizing notions of symmetries to higher-dimensional charged objects [3] have allowed for the extension of the Landau paradigm to describe such phase transitions [4, 5], for a review see [6].
@@
Condensation of charged matter in gauge theories, governed by the Anderson-Higgs mechanism [7, 8, 9, 10], is one of cornerstones in physics. It plays a key role in our understanding of superconductivity phenomenon and clarifies the nature of electroweak interactions within the Standard Model of particle physics. One question which has stood the test of time is the relationship between the Higgs and confined phases of a gauge theory. Let us illustrate this with the 4D compact U(1) gauge theory, described in terms of a U(1) vector potential $A_{\mu}$, i.e. Maxwell theory with magnetic monopoles. In the pure gauge theory at weak coupling there is a deconfined phase in which static electric charges interact via a $1/r$ Coulomb potential mediated by a gapless photon, while at strong coupling the theory enters a confined phase, via the proliferation of magnetic monopoles, where static electric charges interact via a linearly rising potential. If we couple the gauge field to a charged scalar Higgs field, then the Higgs field may condense, leading to a Higgs regime where the photon becomes massive. Can the Higgs regime be sharply distinguished from the confined regime? For a Higgs field in the fundamental representation of the gauge group, i.e. one carrying elementary charge, it is generally understood that the Higgs and confined regimes are actually the same phase, i.e. they are not separated by any thermodynamic bulk phase transition, as shown in [11, 12]. This Higgs-confinement continuity is believed to be true in generic models with a gauge group $\mathcal{G}$ coupled to a scalar Higgs field in the fundamental representation, including both discrete and continuous Abelian and non-Abelian gauge groups [12].
@@
Nonetheless, the question of whether there is a qualitative difference between these two regimes has been revisited time after time from many different perspectives [13]. To survey briefly the history of the endeavor to delineate these two regimes, one approach has been to perform a partial gauge fixing and observe symmetry breaking in an unfixed global subgroup, which shows a phase transition separating them, though the location of the transition line is gauge-dependent and thus lacks a clear physical meaning  [14]. Other proposals seek to delineate them in the presence of global symmetries (whose realization is unaltered between the two regimes), see e.g.  [15, 16, 17]. Yet another approach, advanced partially by one of the authors, emphasizes that (in a certain limit) Abelian Higgs phases with fundamental matter exhibit symmetry-protected topological (SPT) order  [18, 19, 20]. This observation motivates the investigation of the Higgs mechanism in open geometries, and zooms in on low-energy excitations localized near boundaries. In contrast to the confined regime, where the ground state is unique, in the Higgs regime previous studies uncovered energy spectrum degeneracies, see for example [18, 19, 20]. The robustness of these degeneracies originating from boundary-localized modes originates from the interplay of the protecting (generalized) symmetries that depends on the gauge group and dimensionality of the problem. In summary, the presence of a boundary introduces a criterion by which one can delineate the Higgs and confined regimes of Abelian gauge theories with fundamental matter—they are separated by a boundary phase transition.
@@
In this paper we explore in detail boundary symmetry breaking in Wilson lattice gauge theories. We find that the Higgs-confinement boundary criticality mechanism is in fact ubiquitous. We begin in Section II by showing the presence of a boundary phase transition in the $U(1)$ Abelian Higgs model in four dimensions, where the magnetic one-form symmetry is broken explicitly. We discuss how, in the presence of boundaries which allow flux but not charge to exit the system, there is a bulk U(1) global matter symmetry which, by the Gauss law, is equivalent to an electric flux symmetry acting on the boundary. We show that in a particular limit of the theory, in which the action reduces to a 3D XY model on the boundary, this boundary U(1) symmetry can be broken spontaneously. We provide numerical evidence that there is a corresponding boundary phase transition in the presence of bulk fluctuations, and we trace the phase boundary in the bulk phase diagram. Next we turn to non-Abelian Higgs theories in Section III. We consider two types of Higgs models, those with group-valued Higgs fields and those with fundamental representation (vector-valued) Higgs fields, which coincide for gauge group $\mathrm{SU}(2)$ but differ for other gauge groups. We demonstrate that these models also have a global charge symmetry which is realized at the boundary. Using large-scale lattice simulation, we first show that 4D SU(2) Higgs theory has a boundary phase transition in the 3D O(4) universality class, verifying our theoretical prediction. We show that this boundary symmetry breaking is expected to be generic in group-valued Higgs models, and provide a general argument that fundamental-Higgs models with gauge group SU($N$) and SO($N$) exhibit O($2N$) and O($N$) boundary criticalities, respectively. We verify this prediction numerically for the case of the 4D SU(3) fundamental-Higgs. Lastly, in Section IV, we consider generalizing to higher-form Abelian-Higgs models, with a $k$-form gauge field coupled to a $(k-1)$-form Higgs field. We discuss how the higher-form matter symmetry is realized at the boundary through the Gauss law, and show that the action reduces in a limiting case to a boundary $(k-1)$-form gauge theory which may exhibit a confinement-deconfinement phase transition in which the matter symmetry is spontaneously broken. In the same section, we perform a duality transformation and discuss how this symmetry breaking can be viewed from the perspective of magnetic defects which live at the boundary. Finally, we provide an overview of our findings and discuss how our work may connect to symmetry protected topology.
II Boundary Symmetry Breaking in the Abelian Higgs Model
@@
where $\lambda$ is an arbitrary 0-form. By construction, gauge-non-invariant operators cannot exhibit a non-zero vacuum expectation which is formalized by Elitzur’s theorem. Thus, while it is commonly stated that this theory exhibits spontaneous breaking of the U(1) gauge symmetry leading to the Higgs phase, $\langle\phi(x)\rangle=0$ and thus cannot serve as a local order parameter for such a phase transition. Rather the gauge-invariant observables are non-local string operators such as Wilson lines, for example
@@
where $C$ is a curve from $y$ to $x$. When properly normalized, such observables allow one to distinguish quantiatively the deconfined phase from the Higgs-confined phase as discussed in some detail by Fredenhagen and Marcu [21] and [22].
@@
[…]
Figure 1: A sketch of the phase diagram of the 4D U(1) lattice Abelian-Higgs phase diagram discussed in Section II. The confined and Higgs regimes belong to the same thermodynamic phase, though we demonstrate that in presence of symmetry-preserving boundary they are sharply separated by a second order boundary phase transition in the 3D XY universality class. Along the line $\kappa=0$ the model has an exact electric 1-form symmetry, while along the line with $\beta=\infty$ the model has an exact magnetic 1-form symmetry. The confined regime is smoothly connected to the $\beta=0$ limit where the gauge field is maximally disordered. Along the $\kappa=\infty$ line, the bulk is completely frozen, and the boundary reduces to a 3D XY model. The boundary $U(1)$ symmetry is spontaneously broken on the Higgs side of the transition line. A review of the structure of this phase diagram is provided in Appendix B.
II.1 Preliminaries: Lattice Formulation
@@
II.1.1 Action Formulation
To study this theory in more depth we consider regulating it by imposing a UV lattice cutoff. We undertake our exploration of Higgs phases in gauge theories within the Wilson-Fradkin-Shenker lattice formulation. We work on a 4D hypercubic lattice with linear dimension $L$ and periodic boundaries, a discretization of four-dimensional Euclidean spacetime. We consider a a complex 0-form Higgs field taking values $\phi_{i}$ at each vertex $i$. Expanding the Higgs field at site $i$ as $\cramped{\phi_{i}=\rho_{i}\exp(i\theta_{i})}$, we freeze the radial mode by fixing the radius $\rho_{i}$, which does not affect the qualitative physics.111 This freezing corresponds to the limit of infinite bare Higgs self-coupling [12, 23], and the radial mode will be restored upon coarse-graining. Equivalently it may be regarded as a Stückelberg field, and the model can be viewed as a lattice discretization of a gauged nonlinear sigma model with target space U(1). We use these two perspectives to give two different generalizations to non-Abelian gauge groups in Section III. Thus we work with the compact $\mathbbm{R}/2\pi\mathbbm{Z}$-valued 0-form $\theta$, i.e. the phase of the Higgs field, which is minimally coupled to the dynamical U(1) gauge field. We consider a compact 1-form gauge potential $A$ taking values on each oriented link $\ell$, $\cramped{A_{\ell}\in\mathbbm{R}/2\pi\mathbbm{Z}}$. Denoting the reversed orientation by $-\ell$, the gauge potential satisfies $\cramped{A_{-\ell}=-A_{\ell}}$. We will often denote link variables by their endpoints, i.e. $A_{ij}=-A_{ji}$. See Appendix A for a more detailed explanation of lattice differential forms and the notation used here.
We demand the theory to be invariant under gauge transformations of the form of Eq. 2, which on the lattice become
@@ II.1.1 Action Formulation
where $\lambda$ is an arbitrary $\mathbbm{R}/2\pi\mathbbm{Z}$-valued 0-form, and $\mathrm{d}$ is the discrete exterior derivative, $\cramped{(\mathrm{d}\lambda)_{ij}\equiv\lambda_{j}-\lambda_{i}}$. The minimal gauge-invariant building blocks are the Wilson links $\Lambda_{\ell}$, defined on oriented links $\ell$,
@@
which reduces upon substituting in Eqs. 5 and 6 to the Abelian-Higgs model, the lattice equivalent of Eq. 1 in the limit where the radial mode of the Higgs field is frozen,
@@
II.1.2 Hamiltonian Formulation and Gauss Law
It will also serve us to consider the Hamiltonian formulation of the model on a 3D cubic lattice with continuous time. This may be obtained from the action by fixing to temporal gauge ($A_{\ell}=0$ on all timelike links) and taking the continuum limit in the time direction, expressing the partition sum in terms of transfer matrices [24]. The Higgs field phase and gauge connection become operators, denoted $\hat{\theta}_{i}$ and $\hat{A}_{\ell}$ respectively, acting on a local Hilbert space on each vertex or link. They each have a canonically conjugate operator, denoted $\hat{n}_{i}$ and $\hat{E}_{\ell}$ respectively, both with integer eigenvalues, satisfying $[\hat{\theta}_{i},\hat{n}_{i}]=i$ and $[\hat{A}_{\ell},\hat{E}_{\ell}]=i$. Thus $\mathrm{exp}(\pm i\hat{\theta_{i}})$ is the raising/lowering operator for $\hat{n}_{i}$, while $\mathrm{exp}(\pm i\hat{A}_{\ell})$ is the raising/lowering operator for $\hat{E}_{\ell}$. The Hamiltonian may then be expressed as [^2] Note that the $\beta$ and $\kappa$ couplings in the Hamiltonian formulation cannot be quantitatively compared to their values in the Lagrangian formulation, as they are renormalized when taking the continuum limit in the timelike direction [25].
$H_{\text{AH}}=\sum_{\ell}\hat{E}_{\ell}^{2}-{\beta}\sum_{p}\cos(\mathrm{d}\hat{A})_{p}+\sum_{i}\hat{n}_{i}^{2}-{\kappa}\sum_{\ell}\cos(\mathrm{d}\hat{\theta}-\hat{A})_{\ell}.$ (10)
@@ II.1.2 Hamiltonian Formulation and Gauss Law
Gauge transformations, Eq. 4, are implemented by the operators
$\hat{G}[\lambda]=$ $\exp(i\sum_{i}\lambda_{i}\hat{n}_{i}+i\sum_{\ell}(\mathrm{d}\lambda)_{\ell}\hat{E}_{\ell})$
@@
Here we have used the coexterior derivative (see Appendix A)
@@
$-(\mathrm{d}^{\dagger}\hat{E})_{i}=(\nabla\cdot\hat{E})_{i}=\hat{n}_{i},$ (13)
@@
$\hat{W}[\gamma]=\exp\left(i\sum_{\ell\in\gamma}(\mathrm{d}\hat{\theta}-\hat{A})_{\ell}\right),$ (14)
@@
II.1.3 Magnetic Monopoles
In addition to the electric sector, there is also the magnetic sector, though it is not readily seen in this formulation, instead being exposed by duality transformations [26, 27, 28, 29, 30] (see Section IV.2. In the Hamiltonian formulation, the magnetic excitations are sources of divergence of the magnetic field, $\hat{B}=\mathrm{d}\hat{A}$, i.e. magnetic monopoles. In the action formulation, they may be viewed as U(1) vortex defects of the gauge field, characterized by $\mathrm{d}^{2}A\neq 0$, which are allowed because the identity $\mathrm{d}^{2}=0$ is only enforced modulo $2\pi$. In 4D, these homotopy defects form 1-dimensional closed strings in the dual lattice, which we refer to as ’t Hooft loops, the worldlines of magnetic monopoles. They are necessarily included in the Euclidean lattice gauge theory partition sum due to the compactness of the U(1) link variables. In the limit $\beta\to\infty$ of Eq. 8, fluctuations of the gauge field are completely suppressed and no monopoles are present. Correspondingly, we may associate this absence of monopoles to a 1-form symmetry, called magnetic symmetry, indicated in Fig. 1.

### v2

@@ Abstract
[…]
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
The redundancy inherent to gauge theories leads to important subtleties. In particular, while the theories are local, the physical gauge-invariant objects are non-local Wegner-Wilson string loops [1, 2]. This implies a tension in describing the spontaneous breaking of symmetries in the presence of dynamical gauge fields. Whereas many condensed matter systems demonstrate symmetry lowering phase transitions governed by spontaneous breaking of a global symmetry, gauge redundancy, being unphysical, cannot be broken. This fact, enshrined in Elitzur’s theorem [3], belies a rich landscape of different phases separated by phase transitions whose study began systematically in the 1970s—confined, deconfined, Higgs, topologically ordered, etc. While Landau theory successfully accounts for a broad range of phase transitions in correlated many-body systems, the order parameters for theories with dynamical gauge fields are necessarily non-local, raising the question of how to understand the nature of the phase transitions in such theories. Recent advances generalizing notions of symmetries to higher-dimensional charged objects [4] allowed to extend the Landau paradigm to describe such phase transitions [5, 6, 7], for a review see [8].
@@
The Anderson-Higgs mechanism [9, 10, 11, 12], i.e. the condensation of charged scalar matter in gauge theories, is a cornerstone of physics. It plays a key role in our understanding of superconductivity phenomena and the nature of electroweak interactions within the Standard Model of particle physics. Two closely related questions about this mechanism have stood the test of time and continue to generate significant interest: First, what is the gauge-invariant order parameter characterizing Higgs phases [13]? Second, what is the distinction between the Higgs and confined phases [14]? Let us illustrate this with the 4D compact U(1) gauge theory, i.e. Maxwell theory with magnetic monopoles. In the pure gauge theory there is a deconfined phase at weak coupling in which static electric charges interact via a $1/r$ Coulomb potential mediated by a gapless photon. At strong coupling, the proliferation of magnetic monopoles drives the system into the gapped confined phase, where static electric charges interact via a linearly rising potential. On the other hand, if we couple the gauge field to a charged scalar Higgs field, then the Higgs field may condense, driving a transition from the gapless deconfined phase to a gapped Higgs phase via the Anderson-Higgs mechanism. Can these two gapped regimes be distinguished, and if so by what order parameter? For a Higgs field in the fundamental representation of the gauge group, i.e. one carrying elementary charge, it is generally understood that the Higgs and confined regimes are actually the same phase, i.e. they are not separated by any thermodynamic bulk phase transition [15, 14]. This Higgs-confinement continuity is believed to be true in generic models with a gauge group $\mathcal{G}$ coupled to a scalar Higgs field in the fundamental representation, including both discrete and continuous Abelian and non-Abelian gauge groups [14].
@@
The question of whether there is a qualitative difference between these two regimes has been revisited time after time from many different perspectives [16]. To survey briefly the history of the endeavor to delineate these two regimes, one approach has been to perform a partial gauge fixing and observe symmetry breaking in an unfixed global subgroup, which shows a phase transition separating them, though the location of the transition line is gauge-dependent and thus lacks a clear physical meaning  [17]. Other proposals seek to delineate them in the presence of global symmetries (whose realization is unaltered between the two regimes), see e.g.  [18, 19, 20]. Yet another approach, advanced partially by one of the authors, emphasizes that (in a certain limit) Abelian Higgs phases with fundamental matter exhibit symmetry-protected topological (SPT) order  [21, 22, 23]. This observation motivates the investigation of the Higgs mechanism in open geometries, and zooms in on low-energy excitations localized near boundaries. In contrast to the confined regime, where the ground state is unique, in the Higgs regime previous studies uncovered energy spectrum degeneracies, see for example [21, 22, 23]. The robustness of these degeneracies originating from boundary-localized modes arises from the interplay of the protecting (generalized) symmetries that depends on the gauge group and dimensionality of the problem. In summary, the presence of a boundary introduces a criterion by which one can delineate the Higgs and confined regimes of Abelian gauge theories with fundamental matter—they are separated by a boundary phase transition.
@@
In this paper we explore in detail boundary symmetry breaking in Wilson lattice gauge theories. We find that the Higgs-confinement boundary criticality mechanism is in fact ubiquitous. We begin in Section˜2 by showing the presence of a boundary phase transition in the 4D $U(1)$ Abelian Higgs model, where the magnetic one-form symmetry is broken explicitly. We discuss how, in the presence of boundaries which allow flux but not charge to exit the system, there is a bulk U(1) global matter symmetry which, by the Gauss law, is equivalent to an electric flux symmetry acting on the boundary. We show that in a particular limit of the theory, in which the action reduces to a 3D XY model on the boundary, this boundary U(1) symmetry can be broken spontaneously. We provide numerical evidence that there is a corresponding boundary phase transition in the presence of bulk fluctuations, and we trace the phase boundary in the bulk phase diagram. Next we turn to non-Abelian Higgs theories in Section˜3. We consider two types of Higgs models, those with group-valued Higgs fields and those with fundamental representation (vector-valued) Higgs fields, which coincide for gauge group $\mathrm{SU}(2)$ but differ for other gauge groups. We demonstrate that these models also have a global charge symmetry which is realized at the boundary. Using large-scale lattice simulation, we first show that 4D SU(2) Higgs theory has a boundary phase transition in the 3D O(4) universality class, verifying our theoretical prediction. We show that this boundary symmetry breaking is expected to be generic in group-valued Higgs models, and provide a general argument that fundamental-Higgs models with gauge group $\mathrm{SU}(N)$ and $\mathrm{SO}(N)$ exhibit $\mathrm{O}(2N)$ and O($N$) boundary criticalities, respectively. We verify this prediction numerically for the case of the 4D SU(3) fundamental-Higgs. Lastly, in Section˜4, we consider generalizing to higher-form Abelian-Higgs models, with a $k$-form gauge field coupled to a (k1)(k-1)-form Higgs field. We discuss how the higher-form matter symmetry is realized at the boundary through the Gauss law, and show that the action reduces in a limiting case to a boundary (k1)(k-1)-form gauge theory which may exhibit a confinement-deconfinement phase transition in which the matter (k1)(k-1)-form symmetry is spontaneously broken. In the same section, we perform a duality transformation and discuss how this symmetry breaking can be viewed from the perspective of magnetic defects which live at the boundary. Finally, we provide an overview of our findings and discuss how our work gives rise to new insights and challenges for the Higgs=SPT proposal.
[…]
Figure 1: A sketch of the phase diagram of the 4D U(1) lattice Abelian-Higgs phase diagram discussed in Section˜2. The confined and Higgs regimes belong to the same thermodynamic phase, though we demonstrate that in presence of symmetry-preserving boundary they are sharply separated by a second order boundary phase transition in the 3D XY universality class. Along the line $\kappa=0$ the model has an exact electric 1-form symmetry, while along the line with $\beta=\infty$ the model has an exact magnetic 1-form symmetry. The confined regime is smoothly connected to the $\beta=0$ limit where the gauge field is maximally disordered. Along the $\kappa=\infty$ line, the bulk is completely frozen, and the boundary reduces to a 3D XY model. The boundary $U(1)$ symmetry is spontaneously broken on the Higgs side of the transition line. A review of the structure of this phase diagram is provided in Appendix˜B.
2 Boundary Symmetry Breaking in Abelian Higgs Models
@@
where $\lambda$ is an arbitrary 0-form. By construction, gauge-non-invariant operators cannot exhibit a non-zero vacuum expectation which is formalized by Elitzur’s theorem [3]. Thus, while it is commonly stated that this theory exhibits spontaneous breaking of the U(1) gauge symmetry leading to the Higgs phase, $\langle\phi(x)\rangle=0$ and thus cannot serve as a local order parameter for such a phase transition. Rather, the gauge-invariant observables are non-local string operators, such as an open Wilson line
@@
where $C$ is a curve from $y$ to $x$, which creates electric charges attached to their concomitant electric field lines [24, 2]. When properly normalized, such observables allow one to quantitatively distinguish the deconfined phase from the Higgs-confined phase  [25, 26, 27, 28].
@@
2.1 Preliminaries: Lattice Formulation
@@
2.1.1 Action Formulation
To study this theory in more depth we consider regulating it by imposing a UV lattice cutoff. We undertake our exploration of Higgs phases in gauge theories within the Wilson-Fradkin-Shenker lattice formulation. We work on a 4D hypercubic lattice with linear dimension $L$ and periodic boundaries, a discretization of four-dimensional Euclidean spacetime. We consider a a complex 0-form Higgs field taking values $\phi_{i}$ at each vertex $i$. Expanding the Higgs field at site $i$ as $\cramped{\phi_{i}=\rho_{i}\exp(i\theta_{i})}$, we freeze the radial mode by fixing the radius $\rho_{i}$, which does not affect the qualitative physics.111 This freezing corresponds to the limit of infinite bare Higgs self-coupling [14, 26], and the radial mode will be restored upon coarse-graining. Equivalently it may be regarded as a Stückelberg field, and the model can be viewed as a lattice discretization of a gauged nonlinear $\sigma$-model with target space U(1). We use these two perspectives to give two different generalizations to non-Abelian gauge groups in Section 3. Thus we work with the compact $\mathbbm{R}/2\pi\mathbbm{Z}$-valued 0-form $\theta$, i.e. the phase of the Higgs field, which is minimally coupled to the dynamical U(1) gauge field. We consider a compact 1-form gauge potential $A$ taking values on each oriented link $\ell$, $\cramped{A_{\ell}\in\mathbbm{R}/2\pi\mathbbm{Z}}$. Denoting the reversed orientation by -$\ell$, the gauge potential satisfies $\cramped{A_{-\ell}=-A_{\ell}}$. We will often denote link variables by their endpoints, i.e. $A_{ij}=-A_{ji}$. See Appendix˜A for a more detailed explanation of lattice differential forms and the notation used here.
We demand the theory to be invariant under gauge transformations of the form of Eq.˜2, which on the lattice become
@@ 2.1.1 Action Formulation
where $\lambda$ is an arbitrary $\mathbbm{R}/2\pi\mathbbm{Z}$-valued 0-form, and $\mathrm{d}$ is the discrete exterior derivative, defined so that $\cramped{(\mathrm{d}\lambda)_{ij}\equiv\lambda_{j}-\lambda_{i}}$. The minimal gauge-invariant building blocks are the Wilson links $\Lambda_{\ell}$, defined on oriented links $\ell$,
@@
which reduces upon substituting in Eqs.˜5 and 6 to the Abelian-Higgs model, the lattice equivalent of Eq.˜1 in the limit where the radial mode of the Higgs field is frozen,
@@
2.1.2 Hamiltonian Formulation and Gauss Law
[…]
@@ 2.1.2 Hamiltonian Formulation and Gauss Law
Gauge transformations, Eq.˜4, are implemented by the operators
[…]
@@
Here we have used the coexterior derivative (see Appendix˜A)
@@
$(\mathrm{d}^{\dagger}\hat{E})_{i}=(\nabla\cdot\hat{E})_{i}=\hat{n}_{i},$ (13)
@@
$),\hat{W}[\gamma]=\exp\left(i\sum_{\ell\in\gamma}(\mathrm{d}\hat{\theta}-\hat{A}$$_{\ell}\right)$, (14)
@@
2.1.3 Magnetic Monopoles
In addition to the electric sector, there is also the magnetic sector, though it is not readily seen in this formulation, instead being exposed by duality transformations [31, 32, 33, 34, 35] (see Section˜4.2. In the Hamiltonian formulation, the magnetic excitations are sources of divergence of the magnetic field, $\hat{B}=\mathrm{d}\hat{A}$, i.e. magnetic monopoles. In the action formulation, they may be viewed as U(1) vortex defects of the gauge field, characterized by $\mathrm{d}^{2}A\neq 0$, which are allowed because the identity $\mathrm{d}^{2}=0$ is only enforced modulo $2\pi$. In 4D, these homotopy defects form 1-dimensional closed strings in the dual lattice, which we refer to as ’t Hooft loops, the worldlines of magnetic monopoles. They are necessarily included in the Euclidean lattice gauge theory partition sum due to the compactness of the U(1) link variables. In the limit $\beta\to\infty$ of Eq.˜8, fluctuations of the gauge field are completely suppressed and no monopoles are present. Correspondingly, we may equate this absence of monopoles (equivalently, the closure of magnetic flux lines) to the presence of an exact magnetic 1-form symmetry, indicated in Fig.˜1.